OpenStudy (anonymous):

please help:If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

please help me to solve this one

OpenStudy (anonymous):

I think 12/11 of a day

OpenStudy (anonymous):

not so sure

OpenStudy (anonymous):

tell me what is your equation you have?

OpenStudy (amistre64):

how much of the job is done by each person in one day?

OpenStudy (anonymous):

1/2 1/4 and 1/16

OpenStudy (amistre64):

S takes 4 days to do it alone; in 1 day S does 1/4 of the job L takes 6 days to do it alone; in 1 day L does 1/6 of the job T can do it all in 2 days; in 1 day T does 1/2 the job How much is done in one day if they all work together?

OpenStudy (anonymous):

1/2+1/4+1/6

OpenStudy (amistre64):

very good :) 1/2 = 6/12 1/4 = 3/12 1/6 = 2/12 ---------- 11/12 is done in one day.

OpenStudy (amistre64):

since the work is divided into 12 parts; and 11 parts get done each day we get: 11 * days = 12 days = 12/11

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

do you need that in hours and secs and such? or is that a good enough answer :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

haha I was right...

OpenStudy (anonymous):

you great

OpenStudy (amistre64):

:)